OpenStudy (anonymous):

use the distributive property to help you solve the equation 6(w-4)=30 w=?

OpenStudy (anonymous):

you would have 6w - 24=30. you can now solve for w.

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

oops \[\frac{6}{6}\neq0\]

OpenStudy (anonymous):

no. add 24 to both sides giving you 6w=54. Now divide 54/6 to get 9.

OpenStudy (anonymous):

okay so im lost what would the answer be?

OpenStudy (anonymous):

you would get w=9. And this works because when we plug w back in, we would get 6(9-4)=30, which is a true statement.

OpenStudy (anonymous):

aaaah ok